Shift report 20011121 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Wed Nov 21 2001 - 23:56:20 EST

  • Next message: Jens Jørgen Gaardhøje: "BRAHMS intention to submit 200 GeV Mult. Paper announced to RHIC community!"

    ------------------------------------------------------------
    Date: Wed Nov 21 23:56:20 2001, Shift: 16:00-24:00
    Supervisor: rw, with: 
    Log book#11 pages 21 - 22
    
    Shift Summary:
    ------------------------------------------------------------
    1.At 18:20 a good storage was made. Blue/Yellow was 27/32.
    ZDC was 1600. Physics run started at 18:26. The storage was
    ended by QLI at 21.03. Another good storage was made at 22:45.
    blue/yellow=28/28 ZDC rate was 600 and increased to 1000 by 
    steering at 23:30. Physics run started at 22:50.
    
    2.    	run	trigger 6 	setting
    	5931	150K		FS at 4deg.-1/5A, MRS same as the previous
    	5932	150K		same as above
    	5933	26K		FS 12deg-2/3A, MRS same
    	5934	TPC pedestal
    	5935	DVM
    	5936    Pedestals
    	5937	100K		FS MRS same as the previous run
    	5938	running         same as above
    
    3. No new problem has been detected.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Wed Nov 21 2001 - 23:57:10 EST